# Information theory for data-driven model reduction in physics and biology

### v3

@@ Abstract
Model reduction is the construction of simple yet predictive descriptions of the dynamics of many-body systems in terms of a few relevant variables. A prerequisite to model reduction is the identification of these variables, a task for which no general method exists. Here, we develop an approach to identify relevant variables, defined as those most predictive of the future, using the so-called information bottleneck. We elucidate analytically the relation between these relevant variables and the eigenfunctions of the transfer operator describing the dynamics. In the limit of high compression, the relevant variables are directly determined by the slowest-decaying eigenfunctions. Our results provide a firm foundation to interpret deep learning tools that automatically identify reduced variables. Combined with equation learning methods this procedure yields the hidden dynamical rules governing the system’s evolution in a data-driven manner. We illustrate how these tools work in diverse settings including model chaotic and quasiperiodic systems in which we also learn the underlying dynamical equations, uncurated satellite recordings of atmospheric fluid flows, and experimental videos of cyanobacteria colonies in which we discover an emergent synchronization order parameter.
@@
The exhaustive description of a biological or physical system is usually impractical due to the sheer volume of information involved. As an example, the fluids in a steam engine may be described by a $10^{27}$–dimensional state vector containing the positions and momenta of every particle in the engine. Yet, for most practical purposes, its working can be effectively described using only a small number of thermodynamic variables like pressure and temperature. Likewise, some aspects of the behavior of a nematode or a fruit fly, with $385$ and $\sim$140,000 neurons respectively, may be effectively captured by orders of magnitude fewer degrees of freedom [1, 2, 3, 4, 5]. Similar reductions can be achieved for systems ranging from diffusing particles to biochemical molecules and complex networks. In all cases, certain relevant variables can be predicted far into the future even though individual degrees of freedom in the system are effectively unpredictable.
The process by which one goes from the complete description of a system to a simpler one is known as model reduction. Diverse procedures for model reduction exist across the natural sciences. They range from analytical methods, such as adiabatic elimination and multiple-scale analysis [6, 7, 8, 9, 10, 11, 12, 13, 14], to data-driven methods such as dynamic mode decomposition and other linear projection-based methods [15, 16, 17, 18, 19, 20], diffusion maps [21], spectral submanifolds [22, 23], independent component analysis [24], and deep encoder-decoder neural networks [25, 26, 27, 28, 29, 30, 31, 32].
Model reduction consists of two ingredients: a decomposition of the full system into relevant and irrelevant variables, and a self-contained dynamical equation which governs the evolution of the relevant variables. In the absence of prior knowledge and intuition (e.g. a clear separation of scales), identifying a decomposition that produces a minimal set of variables which are maximally descriptive of the full system is an open problem [9]. The answer to this question depends both on the time scale and precision with which one aims to describe the system [33, 34]. In addition, it also depends on one’s ability to identify an adequate dynamical equation for the relevant variables, especially in data-driven settings, where the relevant variables and the dynamical equation are to be learned simultaneously.
To make progress, we develop an information-theoretic framework for model reduction. Very much like MP3 compression is about retaining information that matters most to the human ear [35], model reduction is about keeping information that matters most to predict the future [36, 37]. However, formalizing this intuitive statement requires some care in order to avoid trivial results, that manual approaches usually avoid using physical intuition. By carefully formulating model reduction within this information-theoretic perspective, we show that model reduction (the identification of both relevant variables and their effective dynamics) can be decomposed into a two-step process. First, the problem of variable identification can be reformulated in a way that removes any reference to the effective dynamics, and solved on its own. Second, once the relevant variables are known, an independent dynamical-equation learning step can yield the effective dynamics.
Our key step is to relate the model reduction objective to a lossy compression problem known as the information bottleneck (IB)[38, 39, 36], and to analytically show how and under what conditions the standard operator-theoretic formalism of dynamical systems [40, 27], which underlies most methods of model reduction, naturally emerges from the optimal compression solution. This allows us to give a precise answer to the question of how to identify relevant and irrelevant variables. Crucially, our framework provides a criterion for when to stop increasing the complexity of a minimal model by directly measuring how informative relevant variables are of their own future and of the future of the entire system. Further, it provides a firm foundation to address a practical problem: the construction of deep learning tools to perform model reduction that are guaranteed to be interpretable. This practical realization of our protocol is based on recently-developed deep learning tools that allow scaling our approach to high dimensional systems [41].
The remainder of this paper is organized as follows. Section 1 formalizes model reduction and gives our information-theoretic starting point (see also SI Sections 1-3). Section 2 shows how the optimal information-bottleneck solution identifies the dominant (left) eigenfunctions of the system’s transfer operator (see also SI Section 4), and delineates a procedure to extract the reduced variables. In Sections 3 and 4 we show how information measures capture features of the system’s transfer operator spectrum and how this relates to the question of “when to stop” increasing the complexity of the reduced model. Section 5 (and SI Sections 5-6) shows that our framework can be used to learn relevant variables directly from data using neural networks. We illustrate our approach on several benchmark dynamical systems and demonstrate that it works even on uncurated datasets, e.g. satellite movies of atmospheric flows downloaded from YouTube. SI Section 7 shows applications to chaotic dynamics including the Lorenz-63 system, the Kuramoto-Sivashinsky equation and a hyperchaotic system. In Section 6 we show how our approach, combined with equation learning algorithms, forms a full model reduction pipeline yielding both the reduced variables describing the system and the equations governing its dynamics. Finally, in Section 7 (and SI Section 8) we apply our method to experimental microscopy videos of cyanobacteria colonies, discovering an emergent synchronization order parameter.
I Model reduction as a compression problem
Consider a system whose state is described by a (random) variable $X_{t}$, which might correspond to anything from the position of a single particle to an image of a fluid flow or the fluorescent molecules in a living system (Fig. 1a,b). The full state can be high dimensional, with a number of dimensions equal to the number of particles in a gas, or the number of pixels in an image. It is often difficult to predict the future state $X_{t+\Delta t}$ of such high-dimensional systems due to the enormous computational cost of evolving the system forward in time to a desired accuracy. However, in many cases there exist some functions of the full state $H_{t}=h(X_{t})$ which are low-dimensional and very informative of the future state of the system. These features often correspond to collective variables which evolve slowly in time. While one may guess the appropriate choice of variables $h(X_{t})$ in contexts where all microscopic details of the system are known, we are interested in how to find such variables without such a priori knowledge.
We formulate the task of identifying the reduced variables and their evolution as an optimization problem. A “good” reduced model is one which entails a compressed description of the state, and which may be evolved in time in place of the full system to make predictions. The procedure consists of three steps: first, reduction of the full system by an encoding $h_{\text{enc}}\mathrel{\mathop{\ordinarycolon}}X_{t}\mapsto H_{t}$; second, evolution of the reduced state via an evolution operator $\mathcal{S}\mathrel{\mathop{\ordinarycolon}}H_{t}\mapsto H_{t+\Delta t}$; third, decoding the reduced state via a function $h^{-1}_{\text{dec}}\mathrel{\mathop{\ordinarycolon}}H_{t+\Delta t}\mapsto X_{t}$ to predict the full system’s state. These functions can be found by optimizing
$\min_{h_{\text{enc}},\mathcal{S},h^{-1}_{\text{dec}}}$ $\dim h_{\text{enc}}(X_{t})+\beta\,d\big(h^{-1}_{\text{dec}}(\mathcal{S}h_{\text{enc}}(X_{t})),X_{t+\Delta t}\big).$ (1)
The first term is a compressive constraint enforcing a “reduction” of the model, with the $\dim(\cdot)$ function measuring e.g. the number of bits used to represent the encoded state, or the sparsity of its representation in a preferred basis. The second term captures the error in our prediction of the full state when using the reduced model, measured by a “distance” $d(\cdot,\cdot)$. The parameter $\beta$ sets the trade-off between dimensionality reduction and predictive power. Note that if the encoding $h_{\text{enc}}$ is non-invertible (as is generally the case), the decoder $h^{-1}_{\text{dec}}$ is defined probabilistically via a conditional distribution $h^{-1}_{\text{dec}}(H_{t+\Delta t})\sim p(\cdot|H_{t+\Delta t})$. This is analogous to a procedure producing a thermodynamic ensemble compatible with the state variables. For instance, the canonical ensemble associates to every macrostate $h_{\text{enc}}=(N,V,T)$ defined by the number of particles, the volume, and the temperature a distribution $p_{\text{can}}(\cdot|N,V,T)$ over the set of microstates from which the decoded state is drawn, $h^{-1}_{\text{dec}}\sim p_{\text{can}}(\cdot|N,V,T)$.
@@ I Model reduction as a compression problem
The above objective depends on the choice of distance $d$, which incorporates the geometry of the state space and any desired biases about what is “important” in the system; it may differ from system to system. In what follows, we measure accuracy in terms of the mutual information $I$ shared between the reduced model prediction and the full state. This abstracts away from the geometry of the space and characterizes the similarity of variables in terms of how predictive they are of each other, rather than how close they are, in a system-independent way. This has an important consequence: as a result of the data processing inequality, we can show that our objective function simplifies to one depending only on $h_{\text{enc}}$:
$\min_{h_{\text{enc}}}$ $\quad\dim h_{\text{enc}}(X_{t})-\beta I\big(h_{\text{enc}}(X_{t}),X_{t+\Delta t}\big)$ (2)
@@
The steps to this result are shown in detail in the Methods. This simplification is dramatic: instead of simultaneously learning the encoding function, evolution operator, and decoder, the only step we need to care about is the encoding. Consequently, we view model reduction in the rest of this paper as a task of finding the optimal reduced variables with which to represent our system. The evolution rule $\mathcal{S}$ and decoder $h^{-1}_{\text{dec}}$ are induced by the choice of $h_{\text{enc}}$, and we may view their identification as a secondary, downstream task to be performed after the proper variables have been found.
Before we proceed, we still need to specify the dimensionality reduction. This is subtle, as in different contexts different measures may seem natural, for instance, the entropy for discrete variables, or the number of components for continuous variables. We achieve our dimensionally-reduced variables with a two-step construction. First, we use a term explicitly forcing the reduced representation to lose information about the full state, $\min I(h(X_{t}),X_{t})$ (for notational simplicity, we refer to $h_{\text{enc}}$ as $h$). The minimal objective found in this way
$\min_{h}\mathcal{L}_{\text{IB}}=I(X_{t},h(X_{t}))-\beta I(X_{t+\Delta t},h(X_{t})),$ (3)

### v4

# Information theory for dimensionality reduction in dynamical systems

@@ Abstract
The dynamics of many-body systems can often be captured in terms of only a few relevant variables. Mathematical and numerical approaches exist to identify these variables by exploiting a separation of time scales between slow relevant and fast irrelevant variables, but such a separation of scales is not always obvious or even available. In this work, we introduce an information-theoretic framework for dimensionality reduction in dynamical systems that bypasses this limitation by instead identifying relevant variables based on how predictive they are of the system’s future. To do so, we mathematically formalize the intuition that model reduction is about keeping “relevant” information while throwing away “irrelevant” information. We characterize the solution of the resulting optimization problem and prove that it reduces to standard approaches when a separation of time scales is indeed present in the dynamics. Importantly, we find that within this framework, the problems of identifying relevant variables and identifying their effective dynamics decouple and may be solved separately. This makes the method tractable in practice and enables us to derive dimensionally-reduced variables from data with neural networks. Combined with existing equation learning methods, the procedure introduced in this work reveals the dynamical rules governing the system’s evolution in a data-driven manner. We illustrate these tools in diverse settings including simulated chaotic systems, uncurated satellite recordings of atmospheric fluid flows, and experimental videos of cyanobacteria colonies in which we discover an emergent synchronization order parameter.
@@ I Introduction
The exhaustive description of biological or physical systems is usually impractical due to the sheer volume of information involved. As an example, the fluids in a steam engine may be described by a $10^{27}$–dimensional state vector containing the positions and momenta of every particle in the engine. Yet, for most practical purposes, the working of this engine can be effectively described using only a small number of thermodynamic variables like pressure and temperature, from which most of the observables we care about can be reconstructed. Similar dimensionality reductions can be achieved for dynamical systems spanning the natural sciences.
Standard approaches to dynamical reduction rely on a separation of timescales between slow variables (that we usually care about) and fast variables (that we usually want to remove from the effective description). These approaches, which can be carried out analytically or numerically, include multiple-scale analysis [1], adiabatic elimination [2, 3, 4, 5, 6], center manifold reduction [7, 4], spectral submanifolds [8, 9, 10, 11], renormalization methods [12, 13], diffusion maps [14, 15], or projection-operator (Mori-Zwanzig) methods [16, 17, 18]. These approaches can be augmented by data-driven techniques, some of which have been proposed to occur in the brain [19, 20], ranging from simple statistical methods such as principal component analysis (PCA) to more sophisticated methods like dynamical mode decompositions [21, 22, 23, 24, 25] and neural-network based methods for latent variable learning [23, 26, 27, 28, 29, 30, 31, 32, 33].
However, phenomena that lack a clear timescale separation put these approaches in jeopardy. This lack of scale separation occurs in various situations ranging from climate [34] to ecosystems [35] and disordered systems [36, 37], mainly because many overlapping timescales corresponding to different kinds of interconnected processes are at play. In this situation, we are back to square one: in addition to a scheme that performs model reduction, we need to identify the “latent variables” that are best suited to predict the future of the whole system. Worse, we must solve these two intertwined problems at the same time.
The goal of this work is: (i) to formulate precisely the joint optimization problem that should produce the optimal latent variables and their effective dynamics and (ii) to show that it reduces to the expected result when a separation of timescales is present; and (iii) to implement numerically the optimization procedure in the general case. In order to do so, we develop an information-theoretic framework for collective variable identification (Sec. II). Very much like MP3 compression retains information that matters most to the human ear [38], our reduced description of the system keeps information that matters most to predict the future [39, 40]. Formalizing this intuitive statement requires some care in order to avoid trivial results that manual approaches usually avoid through physical intuition.
A key result of our work is that, within this information-theoretic framework, the identification of the latent variables and of their effective dynamics can be performed separately (Sec. II.2.4). The corresponding optimization problem can then be related to a lossy compression problem known as the information bottleneck [41, 42, 39] (Sec. II.2.5). Once the latent variables are identified, existing methods can readily yield the effective dynamics [43, 44] (examples are carried out in detail in Sec. IV).
We show analytically how and under what conditions the standard operator-theoretic formalism of dynamical systems, which underlies most methods of model reduction, naturally emerges from the optimal compression solution (Sec. III). In particular, we show analytically that our method yields the dominant eigenfunction of the evolution (transfer) operator in limit of high compression under the assumption of Markovian dynamics. As a consequence, our framework reproduces the expected result when a separation of timescales is present, namely that the latent variables are the slow variables. In the context of deep learning, this result also sheds light on the workings of the variational autoencoder-like neural networks that solve the variational information bottleneck problem [45, 46, 47].
In addition, our information-theoretic framework provides criteria for when to stop increasing the complexity of the latent description, by directly measuring how informative relevant variables are of their own future and of the future of the entire system (Sec. IV.1). We also discuss how to quantify and possibly reduce the non-Markovianity of latent variables (Sec. IV.2). Finally, we illustrate how to combine our approach with equation learning algorithms to produce a fully-fledged model discovery pipeline yielding both the latent variables and their effective dynamics (Sec. IV.3).
Section V showcases our framework in uncontrolled situations: satellite movies of atmospheric flows downloaded from YouTube (Sec. V.1) and experimental measurements of experimental microscopy videos of cyanobacteria colonies (Sec. V.2). In these situations, multiple timescales that are not necessarily well-separated are present (like circadian rhythms and cell growth in cyanobacteria experiments, Sec. V.2), the Markovian hypothesis used in Sec. III is not guaranteed to hold, and the data is high-dimensional, precluding naive optimization procedures. Nevertheless, our procedure produces meaningful interpretable latent variables that can be related to the underlying physics or biology.
II Model discovery as compression
II.1 Dimensionality reduction for dynamics
Consider a system whose state is described by a (possibly random) variable $X_{t}$, which might correspond to anything from the position of a single particle to an image of a fluid flow or the fluorescent molecules in a living system (Fig. 1a,b). The full state can be high dimensional, with a number of dimensions equal to the number of particles in a gas, or the number of pixels in an image. The system’s dynamics are described by an evolution operator $\mathcal{U}_{\Delta t}$ that produces the future state of the system $X_{t+\Delta t}=\mathcal{U}_{\Delta t}(X_{t})$ as a function of the current state $X_{t}$, represented graphically as:
$X_{t}$$X_{t+\Delta t}$physical dynamics $\mathcal{U}_{\Delta t}$[Uncaptioned image] (1)
Even when $X_{t}$ is very high dimensional, in many cases it turns out that these dynamics are effectively low-dimensional. This means that we can compress $X_{t}$ into a low-dimensional “latent” variable $H_{t}$ using an encoding function $h$:
$X_{t}$$H_{t}$$h(X_{t})$encoder $h$ (2)
where we have represented the encoder $h$ with a funnel to convey the idea that it should compress information. The dynamics of the latent variable are described by an operator $\mathcal{V}_{\Delta t}$ that acts on the reduced state $H_{t+\Delta t}=\mathcal{V}_{\Delta t}(H_{t})$:
$H_{t}$$H_{t+\Delta t}$latent dynamics $\mathcal{V}_{\Delta t}$[Uncaptioned image] (3)
We can then attempt to recover the future state of the full system by inverting the encoding with a decoding function $\tilde{h}^{-1}$ acting as
$\tilde{X}_{t+\Delta t}$$\equiv\tilde{h}^{-1}(H_{t+\Delta t})$$H_{t+\Delta t}$decoder $\tilde{h}^{-1}$ (4)
where we have represented the decoder as a inverted funnel. In general, information is lost during the encoding procedure and as a result the predicted state and the true state will differ. Consequently, there is finite distance
$d=\lVert X_{t+\Delta t}-\tilde{h}^{-1}(H_{t+\Delta t})\rVert$ (5)
that we want to make as small as possible. In other words, we want to minimize the non-commutativity of the diagram
[…]
However, minimizing $d$ is not our only goal. If it were so, we could simply set $h$ and $\tilde{h}^{-1}$ to the identity and $\mathcal{V}=\mathcal{U}$ to perfectly reproduce the original dynamics. Instead, we wish to simultaneously minimize $d$ while ensuring that the size of the latent variable $H_{t}$ (e.g. its dimension) is as small as possible. To find a good latent representation of the state and its dynamics, we encode these requirements in an optimization problem
[…]
so that the optimal encoder, latent evolution, and decoder (decorated with a star) minimize the objective function (loss function) $\mathcal{L}=\text{size}(H_{t})+\beta\,{d}$, where $\beta$ determines the trade off between predictive accuracy and dimensionality reduction.
The optimization problem Eq. (7) is daunting because it requires simultaneously optimizing over three complex objects. This can be realized in certain settings [48, 49], but it is too difficult in general. It turns out that this difficulty is drastically alleviated by a careful implementation of the idea that we now sketch. Up to now, we have glossed over two key points. First, the encoder is in general not invertible, which naturally leads to a probabilistic decoder. Second, the distance function used to compute $d$ must be chosen to be consistent with this probabilistic picture. As we shall see, a careful choice of distance dramatically simplifies the optimization problem.
II.2 Information theory for dimensionality reduction
II.2.1 Compression requires noninvertible encoders
As $X_{t}$ is a very-high dimensional variable while $H_{t}$ is (hopefully) a low-dimensional one, the encoding $h$ will generally be non-invertible. The non-invertibility of the encoder is the essence of the compression: it is a feature, not a bug. The requirement that $H_{t}$ does not contain “too much” information is contained in the quantity $\text{size}(H_{t})$ that we try to make as small as possible in Eq. (7).
The precise meaning of $\text{size}(H)$ depends on the nature of the variable $H$. When $H$ lives in a discrete space (for instance, if it is a finite chain of bits), then the cardinality of the set of possible states is a reasonable measure of $\text{size}(H)$. When $H$ is a continuous variable (say $H\in\mathbb{R}^{n}$), the cardinality of the set of possible $H$ is infinite, and is therefore not a good measurement of the available size. In principle, one can encode an arbitrarily large amount of information in just a single real number (e.g. in all the decimals). However, the dimension of a space is preserved by continuous bijections [50]. This suggests that in this case, the dimension of (the space where lives) $H_{t}$ is an reasonable measure of $\text{size}(H)$, provided that we impose that the encoder is continuous. The addition of noise into $h$ also limits how much information about $X_{t}$ may be stored in $H_{t}$ [51, 52]. In this case, the size of $H_{t}$ may be captured by e.g. the entropy of the distribution $p(H_{t})$.
II.2.2 Noninvertible encoders require probabilistic decoders
The “inverse” of a non-invertible encoder could produce any of the possibly many states $X_{t}$ that map to a given value of $H_{t}$. The decoder $\tilde{h}^{-1}$ is therefore defined probabilistically via a conditional distribution: instead of obtaining $\tilde{X}_{t+\Delta t}$ as $\tilde{h}^{-1}(H_{t+\Delta t})$, it is a random variable with distribution $p(\tilde{X}_{t+\Delta t}|H_{t+\Delta t})$ and we can think of $\tilde{h}^{-1}(H_{t+\Delta t})$ as a sample from this distribution. Hence, Eq. (4) can more precisely be represented as
[Uncaptioned image] (8)
This is analogous to a procedure producing a thermodynamic ensemble compatible with the state variables. For instance, the canonical ensemble associates to every macrostate $H_{t}=(N,V,T)$ defined by a number of particles, volume, and temperature a distribution $p_{\text{can}}(\xi|N,V,T)$ over the set of microstates $\xi$ from which the decoded state is drawn, namely, $\xi=\tilde{h}^{-1}(H_{t})$ follows the distribution $p_{\text{can}}(\xi|N,V,T)$.
II.2.3 Noise makes dynamics fathomable
The necessity of a probabilistic decoder leads us to embrace the probabilistic nature of all quantities involved: indeed, $X_{t}$, $H_{t}$, and so on can all be treated on an equal footing as random variables  [^1] From now on, we denote random variables by uppercase $X_{t}$ and their values by lowercase $x_{t}$., which allows us to handle stochastic (noisy) evolutions. This turns out to be a key element in our analysis. The idea that randomness regularizes dynamics has emerged in the study of many-body dynamical systems, where chaos provides a justification for the introduction of statistical ensembles [54, 55]. This idea is key in reconciling deterministic dynamics with thermodynamic and hydrodynamic approaches. Indeed, noise (in the form of internal chaotic dynamics and/or external perturbation) tends to induce dynamics that are effectively low-dimensional on long time scales [56, 57, 58]. From a mathematical perspective, this can be understood in terms of the spectrum of the transfer operator associated to the physical dynamics $\mathcal{U}_{\Delta t}$, which may, in certain cases, be split into a finite-dimensional part corresponding to the slow dynamics, plus a (possibly infinite-dimensional and complicated) part that can be neglected on long time scales because it describes the fast dynamics. This is illustrated in Fig. 1d, see Sec. III for technical details. This idea has forcefully been highlighted in Ref. [56], where it is argued that noise limits how much we can resolve states in a chaotic limit, suggesting the existence of a finite-dimensional description whose dimension is limited by the dynamics rather than our observational abilities. Mathematically, this translates into the existence of an optimal partition of phase space associated to a finite-dimensional probabilistic description [59, 57, 56]. Our information-theoretic perspective allows us to extend these ideas to the case of continuous variables, as we discuss in detail in Sec. IV.1.
II.2.4 Information theoretic distance separates concerns
The distance $d$ in Eq. (5) should be compatible with the probabilistic nature of the decoder, and of the other quantities. In particular, $d$ should compare the statistics of decoded states $\tilde{X}_{t+\Delta t}$ with those of the true $X_{t+\Delta t}$, rather than comparing individual samples. To do so, we measure how predictive the latent variable $H_{t+\Delta t}$ is of the future state $X_{t+\Delta t}$ in terms of the mutual information between these variables (SI Sections S1 and S2) by setting
$d=-I\left([\tilde{h}^{-1}\circ\mathcal{V}_{\Delta t}\circ h](X_{t}),X_{t+\Delta t}\right).$ (9)
As shown in Appendix A, mutual information emerges naturally when taking into account the probabilistic nature of the decoder. Notice the sign on the right-hand side of the equation: to minimize the distance, we should maximize the mutual information.
@@ II.2.4 Information theoretic distance separates concerns
The mutual information is not a true metric function, which in turn allows us to abstract away from the geometry of the underlying space. Instead of measuring the similarity of variables in terms of how close they are, it is measured in terms of how predictive they are of each other in a system-independent way. This has an important consequence: as a result of the data processing inequality, we can show that our objective function then simplifies to one depending only on $h$. More precisely, we prove in Appendix B that the optimal encoder is given by
[…]
This simplification is dramatic: instead of simultaneously learning the encoding function , the latent evolution operator [Uncaptioned image], and the decoder , we only need to care about the encoding.
@@
In particular, we can completely forget about the latent evolution at this stage. In the language of computer science, this implements a separation of concerns [60] between the choice of optimal latent variables and the discovery of an optimal latent evolution. Consequently, we view model discovery in the rest of this paper as a task of finding the optimal reduced variables with which to represent our system. The evolution rule $\mathcal{V}$ and decoder $\tilde{h}^{-1}$ are induced by the choice of $h$, and we may view their identification as a secondary, downstream task to be performed after the proper variables have been found.
II.2.5 Continuous information-theoretic sizes help optimization
In Sec. II.2.1, we have focused on situations where the $\text{size}(H_{t})$ is a discrete quantity, like a number of bits or a dimension. However, it is not practical to have discrete quantities in optimization problems. Using an information-theoretic constraint, we can limit the effective size available to $H_{t}$ without enforcing a hard constraint. To do so, we replace in the cost function $\text{size}(H_{t})$ by the term $I(H_{t},X_{t})$ that explicitly forces the reduced representation to lose information about the full state. The objective found in this way
$\mathcal{L}_{\text{IB}}=I(H_{t},X_{t})-\beta I(H_{t},X_{t+\Delta t}),$ (11)